# Knowledge Graph Supported Benchmark and Video Captioning for Basketball

### 25 Jan 2024 v1

@@ Abstract
Despite the recent emergence of video captioning models, how to generate the text description with specific entity names and fine-grained actions is far from being solved, which however has great applications such as basketball live text broadcast. In this paper, a new multimodal knowledge supported basketball benchmark for video captioning is proposed. Specifically, we construct a Multimodal Basketball Game Knowledge Graph (MbgKG) to provide knowledge beyond videos. Then, a Multimodal Basketball Game Video Captioning (MbgVC) dataset that contains 9 types of fine-grained shooting events and 286 players’ knowledge (i.e., images and names) is constructed based on MbgKG. We develop a novel framework in the encoder-decoder form named Entity-Aware Captioner (EAC) for basketball live text broadcast. The temporal information in video is encoded by introducing the bi-directional GRU (Bi-GRU) module. And the multi-head self-attention module is utilized to model the relationships among the players and select the key players. Besides, we propose a new performance evaluation metric named Game Description Score (GDS), which measures not only the linguistic performance but also the accuracy of the names prediction. Extensive experiments on MbgVC dataset demonstrate that EAC effectively leverages external knowledge and outperforms advanced video captioning models. The proposed benchmark and corresponding codes will be publicly available soon.
1 Introduction
Video captioning (VC) is a crucial computer vision task that requires the model to outputs corresponding text descriptions based on the given video. Despite the recent rapid development of video captioning [17, 12, 16, 13, 36, 33], how to automatically generate a text description with specific entity names and fine-grained actions is seldom investigated. As the basketball live text broadcast example in Fig. 1, the basketball video involves multi-person actions and complex scenes, which pose significant challenges to model’s performance and generalization. Note that conventional models can only generate the simple sentence to describe the video from a macroscopic perspective (e.g., a man fails to make a shot and another man gets the rebound). In contrast, if the model has game-related knowledge such as players who appear in the game and fine-grained actions in basketball, it can generate the knowledge-grounded text description (e.g., Brandon Ingram misses the 2pt jump shot and Justise Winslow gets the defensive rebound).
@@ 1 Introduction
In addition, existing common used benchmarks, including MSVD [2], YouCook [3], MSR-VTT [31], and ActivityNet Captions [9], simplify the task by using indefinite pronouns like “a man”, “a woman” or “a group of men” instead of specific entity names. And the actions in the annotated captions are coarse-grained. These benchmarks cannot provide relevant knowledge beyond videos and fine-grained action annotations to develop models to generate text descriptions with specific entity names and fine-grained actions. To our best knowledge, there is no relevant benchmark in basketball that can support this task.
@@
To narrow this gap, a new multimodal knowledge graph supported basketball benchmark for video captioning is proposed. This benchmark requires the model to comprehend the video’s content and generate event descriptions that incorporate specific entities with additional knowledge. Specifically, we collect 25 NBA full-games multimodal data from professional basketball platforms in 2022-2023 season, including events, players’ information, teams’ information, and videos. Based on the collected data, we construct a Multimodal Basketball Game Knowledge Graph (MbgKG) to provide the knowledge beyond videos, as depicted in Fig. 2. Subsequently, a Multimodal Basketball Game Video Captioning (MbgVC) dataset is constructed based on nodes in MbgKG and relationships among the selected nodes. MbgVC dataset comprises 9 types of basketball shooting events and 286 players’ knowledge (i.e., images and names), with data samples illustrated in Fig. 3.
To effectively utilize knowledge in the MbgVC dataset, we propose an encoder-decoder framework named Entity-Aware Captioner (EAC) for basketball live text broadcast. EAC is comprised of 3 separate unimodal encoders for videos, players’ images, and players’ names, as well as a pre-trained language model that serves as the text decoder for generating descriptions of the given videos. Moreover, the Bi-GRU module is introduced to encode the temporal information, while the multi-head self-attention module is introduced to model the associations among the candidate players and select the key players. In addition, we propose a new performance evaluation metric, the Game Description Score (GDS), which takes into account both the linguistic quality of the description (as used in conventional video captioning tasks) and the accuracy of name predictions.
@@
  • A Multimodal Basketball Game Knowledge Graph (MbgKG) is constructed to provide additional contextual knowledge that extends beyond the video contents.
  • Based on MbgKG, nodes and relationships are extracted to construct our Multimodal Basketball Game Video Captioning (MbgVC) benchmark. Additionally, we propose a new evaluation metric for MbgVC that offers a more accurate and comprehensive evaluation by incorporating the accuracy of name predictions.
@@
  • We develop an Entity-aware Video Captioning framework in encoder-decoder form, which selectly incorporates key information from external knowledge to generate text descriptions with specific entity names.
@@
  • On the MbgVC, our proposed model outperforms existing advanced models, achieving leading performance.
@@
Figure 3: Data samples from the proposed dataset. Each video is annotated by fileid, action type, caption, player images and player names. Each of the players involved in caption as well as their teammates serve as candidate players.
2 Related Works
2.1 Computer Vision with Knowledge Graph
@@
A knowledge graph (KG) is essentially a large-scale semantic network that contains entities, concepts as nodes and various semantic relationships among them as edges [37]. The application of knowledge graphs have greatly promoted the rapid development of computer vision. Zhuo et al. [38] utilize the mined category-attribute relationships in knowledge graph and the similarity between the seen categories and the unseen ones for more reliable knowledge transfer. One of the main challenges in video-text retrieval task is to identify fine-grained semantic associations between video and text. To address this issue, under the guidance of external knowledge, Fang et al. [4] utilize the associations between concepts to extend new concepts and enrich the representation of videos. This approach enables a more accurate matching of video and text, leading to improved retrieval results. Gu et al. [6] propose a text with knowledge graph augmented transformer, which integrates external knowledge and leverages multimodal information to mitigate the challenges posed by long-tail words. However, unlike the aforementioned works that extract knowledge from unimodal knowledge graphs, our approach extracts knowledge from a multimodal knowledge graph, utilizing this multimodal knowledge to assist the model in generating descriptions that include specific entity names and fine-grained actions.
2.2 Video Captioning Benchmarks
Existing widely used benchmarks to support the video captioning task include MSVD [2], MSR-VTT [31], YouTube Hightlight [25] and VATEX [30]. These benchmarks are open-domain databases that enhance the generalization of models. However, these benchmarks suffer from common limitations: the text descriptions are too concise, ignoring the names of specific entities and fine-grained action types. Some works also attempt to enrich benchmarks and develop models to generate text descriptions with more fine-grained information. Fang et al. [5] construct a benchmark annotated with captions and commonsense descriptions. This commonsense benchmark develops models to generate captions as well as 3 types of commonsense descriptions (intention, effect, and attribute). Yu et al. [34] propose a fine-grained sports narrative benchmark that focuses more on the detailed actions of the subjects, but this benchmark ignores the names of specific subjects. To assist visually impaired individuals in enjoying movies, Yue et al. [35] construct a large-scale Chinese movie benchmark, which requires models to generate role-aware narration paragraphs when there are no actors speaking. Qi et al. [19] utilized a unimodal knowledge graph for real-time soccer commentary, incorporating all pre-match and player-team information without a selection mechanism. This indiscriminate inclusion could introduce noise, affecting caption accuracy. Contrastingly, our approach first employs a multimodal knowledge graph, integrating both textual and visual information, such as player names and images. This enriches the context for video captioning, improving the match with video content. Second, we propose a selective attention mechanism, modeling associations among candidate players and incorporating key players’ knowledge. This effectively reduces noise and enhances caption precision. These advancements distinguish our work from Qi et al.’s [19] and other existing models, demonstrating the potential of multimodal knowledge and selective attention in video captioning tasks.
2.3 Video Captioning Models
Figure 4: Illustration of the proposed video captioning model that leverages external knowledge from multimodal knowledge graph.
@@ 2.3 Video Captioning Models
Video captioning is a crucial task in video understanding, where models generate text descriptions for given videos. Some works [10, 14, 16, 36] utilize a visual encoder to extract video representations from a set of video frames, and a language decoder transfer these representations to the corresponding descriptions. Recently, CLIP (Contrastive Language-Image Pretraining) [20] has demonstrated its superior performance on various visual-linguistic tasks relying on large-scale contrastive pre-training with image-text pairs. Therefore, Clip4Caption [26] employs CLIP to acquire the aligned visual-text representation, leading to significant improvements in video captioning performance. Benefiting from the flexibility of the transformer [27] architecture, SwinBert [12] introduces Video Swin Transformer [15] as the video encoder to encode spatial-temporal representations from video frames. However, none of the previously mentioned works can address the challenge of generating action entity names effectively. In stark contrast, we propose to incorporate external knowledge based on an encoder-decoder model structure to generate text descriptions with specific entity names and fine-grained actions for basketball live text broadcasts, demonstrating the practicality of our entity-aware video captioning in real applications.
3 Proposed Benchmark
@@
In this section, we first construct a multimodal knowledge graph as the knowledge base, and then extract relevant data and knowledge to construct a multimodal basketball dataset. Furthermore, a new metric for this dataset is proposed to assess the performance of models.
3.1 Multimodal Basketball Game Knowledge Graph
@@
To structure the collected data, the following steps are needed to process it: (1) Categorize basketball events. (2) Parse and structure the descriptions of each type of event. (3) Match these event descriptions with the corresponding video timestamps using OCR toolkits.

### 28 Feb 2024 v2

# Knowledge Guided Entity-aware Video Captioning and A Basketball Benchmark

@@ Abstract
Despite the recent emergence of video captioning models, how to generate the text description with specific entity names and fine-grained actions is far from being solved, which however has great applications such as basketball live text broadcast. In this paper, a new multimodal knowledge graph supported basketball benchmark for video captioning is proposed. Specifically, we construct a multimodal basketball game knowledge graph (KG_NBA_2022) to provide additional knowledge beyond videos. Then, a multimodal basketball game video captioning (VC_NBA_2022) dataset that contains 9 types of fine-grained shooting events and 286 players’ knowledge (i.e., images and names) is constructed based on KG_NBA_2022. We develop a knowledge guided entity-aware video captioning network (KEANet) based on a candidate player list in encoder-decoder form for basketball live text broadcast. The temporal contextual information in video is encoded by introducing the bi-directional GRU (Bi-GRU) module. And the entity-aware module is designed to model the relationships among the players and highlight the key players. Extensive experiments on multiple sports benchmarks demonstrate that KEANet effectively leverages extera knowledge and outperforms advanced video captioning models. The proposed dataset and corresponding codes will be publicly available soon.
Index Terms:
Knowledge Graph, Video captioning, Entity-aware, Basketball.
I Introduction
Video captioning (VC) is a crucial computer vision task that requires the model to outputs corresponding text descriptions based on the given video. By associating visual information and text elements, this task has been blossoming and gaining increasing attention owing to its many promising applications, such as video title automatic generation [1], visually-impaired assistance [2], video storytelling [3] and online video search [4, 5].
Despite the recent rapid development of video captioning [6, 7, 8, 9, 10, 11, 12], these methods are not practical in real-world scenarios. They cannot automatically generate a text description with specific entity names and fine-grained actions. As the basketball live text broadcast example in Fig. 1, the basketball video involves multi-person actions and complex scenes, which pose significant challenges to model’s performance and generalization. Note that conventional models can only generate the simple sentence to describe the video from a macroscopic perspective (e.g., a man fails to make a shot and another man gets the rebound). In contrast, if the model has game-related knowledge such as players who appear in the game and fine-grained actions in basketball, it can generate the knowledge-grounded text description (e.g., Brandon Ingram misses the 2pt jump shot and Justise Winslow gets the defensive rebound). In addition, existing common used benchmarks, including MSVD [13], YouCook [14], MSR-VTT [15], and ActivityNet Captions [16], simplify the task by using indefinite pronouns like “a man”, “a woman” or “a group of men” instead of specific entity names. And the actions in the annotated captions are coarse-grained. These benchmarks cannot provide relevant knowledge beyond videos and fine-grained action annotations to develop models to generate text descriptions with specific entity names and fine-grained actions.
@@ I Introduction
An increasing number of researchers have discovered that conventional methods and benchmarks for video captioning fail to meet the requirements of practical applications and have made numerous attempts to address this issue. Mkhallati et al. [17] publicly release SoccerNet-Caption, the first dataset for dense video captioning in soccer broadcast videos. Although this work focuses on researching entity-aware video captioning, it cannot provide any additional knowledge for the model to generate text descriptions with specific entity names. Ayyubi et al. [18] first train an entity perception detector to detect entities in video frames. They then utilize a large language model to retrieve relevant knowledge and enrich the visual content, thereby generating entity-aware text descriptions for news videos summary. However, this approach is limited by the performance of entity perception detector and large language model. Furthermore, Qi et al. [19] build a unimodal knowledge graph and a benchmark (Goal) for soccer commentary. The provided knowledge is not comprehensive and lacks visual aspects. It can be seen that these methods are tailored to different domains. The introduction of relevant background knowledge is essential to achieve text descriptions with specific entity names in a particular field.
@@
In this work, we focus on the task of knowledge guided entity-aware basketball video captioning (KEBVC), which requires the model to comprehend the content of the provided video and generate text descriptions that include specific entity names and fine-grained actions based on the additional knowledge. We propose a new multimodal knowledge graph supported video captioning benchmark for basketball live text broadcast. This benchmark requires the model to comprehend the video’s content and generate event descriptions that incorporate specific entities with additional knowledge. Specifically, we collect 25 NBA full-games multimodal data from professional basketball platforms in 2022-2023 season, including events, players’ information, teams’ information, and videos. Based on the collected data, we construct a multimodal basketball game knowledge graph named KG_NBA_2022 to provide the knowledge beyond videos, as depicted in Fig. 2. Subsequently, a multimodal basketball game video captioning dataset named VC_NBA_2022 is constructed based on nodes in KG_NBA_2022 and relationships among the selected nodes. VC_NBA_2022 dataset comprises 9 types of basketball shooting events and 286 players’ knowledge (i.e., images and names), with data samples illustrated in Fig. 3.
To effectively utilize knowledge in the VC_NBA_2022 dataset, we propose a knowledge guided entity-aware network (KEANet) based on a candidate player list in encoder-decoder form for basketball live text broadcast. KEANet is comprised of 3 separate unimodal encoders for videos, players’ images, and players’ names, as well as a pre-trained language model that serves as the text decoder for generating descriptions of the given videos. Moreover, the Bi-GRU module is introduced to encode the temporal contextual information, while the entity-aware module is designed to model the associations among the candidate players and highlight the key players.
@@
  • We provide an in-depth analysis to discover characteristics of basketball domain and construct a multimodal Basketball Game Knowledge Graph. This knowledge graph introduce additional knowledge that extends beyond the video contents. Based on this Knowledge Graph, nodes and relationships are extracted to construct our Multimodal Basketball Game Video Captioning benchmark.
@@
  • We develop a knowledge guided entity-aware video captioning network based on a candidate player list in encoder-decoder form, which attentively incorporates key information from additional knowledge to generate text descriptions with specific entity names.
@@
  • To validate the generalization of the proposed model, we conduct experiments on the proposed basketball dataset and the football domain dataset Goal [19]. The proposed model outperforms existing advanced models, achieving leading performance.
@@
Figure 3: Data sample from the proposed dataset. Each video is annotated by fileid, action type, caption, player images and player names. Each of the players involved in caption as well as their teammates serve as candidate players.
II Related Works
II-A Computer Vision with Knowledge Graph
@@
A knowledge graph (KG) is essentially a large-scale semantic network that contains entities, concepts as nodes and various semantic relationships among them as edges [20]. The application of knowledge graphs have greatly promoted the rapid development of computer vision. Zhuo et al. [21] utilize the mined category-attribute relationships in knowledge graph and the similarity between the seen categories and the unseen ones for more reliable knowledge transfer. One of the main challenges in video-text retrieval task is to identify fine-grained semantic associations between video and text. To address this issue, under the guidance of additional knowledge, Fang et al. [22] utilize the associations between concepts to extend new concepts and enrich the representation of videos. This approach enables a more accurate matching of video and text, leading to improved retrieval results. Gu et al. [23] propose a text with knowledge graph augmented transformer, which integrates additional knowledge and leverages multimodal information to mitigate the challenges posed by long-tail words.
However, unlike the aforementioned works that extract knowledge from unimodal knowledge graphs, our approach extracts knowledge from a multimodal knowledge graph, utilizing this multimodal knowledge to assist the model in generating descriptions that include specific entity names and fine-grained actions.
II-B Video Captioning
@@
Video captioning is a crucial task in video understanding, where models generate text descriptions for given videos. Some works [24, 25, 8, 26, 27, 28, 29, 30] utilize a visual encoder to extract video representations from a set of video frames, and a language decoder transfer these representations to the corresponding descriptions. Recently, CLIP (Contrastive Language-Image Pretraining) [31] has demonstrated its superior performance on various visual-linguistic tasks relying on large-scale contrastive pre-training with image-text pairs. Therefore, Clip4Caption [32] employs CLIP to acquire the aligned visual-text representation, leading to significant improvements in video captioning performance. Benefiting from the flexibility of the transformer [33] architecture, SwinBert [7] introduces Video Swin Transformer [34] as the video encoder to encode spatial-temporal representations from video frames. None of the previously mentioned works can address the challenge of generating action entity names effectively. In stark contrast, we propose to incorporate additional knowledge based on an encoder-decoder model structure to generate text descriptions with specific entity names and fine-grained actions for basketball live text broadcasts, demonstrating the practicality of our entity-aware video captioning in real applications.
II-C Video Captioning Benchmarks
Existing widely used benchmarks to support the video captioning task include MSVD [13], MSR-VTT [15], YouTube Hightlight [35] and VATEX [36]. These benchmarks are open-domain databases that enhance the generalization of models. However, these benchmarks suffer from common limitations: the text descriptions are too concise, ignoring the names of specific entities and fine-grained action types. Some works also attempt to enrich benchmarks and develop models to generate text descriptions with more fine-grained information. Fang et al. [37] construct a benchmark annotated with captions and commonsense descriptions. This commonsense benchmark develops models to generate captions as well as 3 types of commonsense descriptions (intention, effect, and attribute). Yu et al. [38] propose a fine-grained sports narrative benchmark that focuses more on the detailed actions of the subjects, but this benchmark ignores the names of specific subjects. To assist visually impaired individuals in enjoying movies, Yue et al. [39] construct a large-scale Chinese movie benchmark, which requires models to generate role-aware narration paragraphs when there are no actors speaking. Byeong et al. [40] propose a benchmark for automatically generating commentary on baseball games. The descriptions in this benchmark focus on player categories rather than specific player names. To achieve meaningful news summarization, Ayyubi et al. [18] propose the task of summarizing news video directly to entity-aware captions. They also release a large-scale dataset, to support research on this task. Mkhallati et al. [17] release a benchmark for soccer broadcasts commentaries. Although the text caption in the benchmark contains specific player names, they do not provide any experimental support for generating a caption with specific names. On the contrary, each specific name is replaced by a specific token (e.g., [TEAM], [COACH], [REFEREE], and [PLAYER]).
Figure 4: Illustration of extracting relevant data using relationship extraction from the knowledge graph and constructing the dataset.
Qi et al. [19] utilized a unimodal knowledge graph for real-time soccer commentary, incorporating all pre-match and player-team information without a selection mechanism. This indiscriminate inclusion could introduce noise, affecting caption accuracy. Contrastingly, our approach first employs a multimodal knowledge graph, integrating both text and visual information, such as player names and images. This enriches the context for video captioning, improving the match with video content. Second, we propose a entity-aware module, modeling associations among candidate players and incorporating key players’ knowledge. This effectively reduces noise and enhances caption precision. These advancements distinguish our work from Qi et al.’s [19] and other existing models, demonstrating the potential of multimodal knowledge and entity-aware in video captioning tasks.
III Proposed Benchmark
@@
In this section, we first construct a multimodal knowledge graph as the knowledge base, and then extract relevant data and knowledge to construct a multimodal basketball dataset.
III-A Multimodal Basketball Game Knowledge Graph
@@
To structure the collected data, the following steps are needed to process it: (1) Categorize basketball events. (2) Parse and structure the descriptions of each type of event. (3) Match these event descriptions with the corresponding video timestamps.
By analyzing the play-by-play data, game events can be divided into 9 categories, including “Foul”, “Rebound”, “Violation”, “Timeout”, “Freethrow”, “Enter game”, “Turnover”, “Jump ball”, and “Shot”. Each text description has its owns keyword, such as the description of the “Foul” event “Personal foul by G. Temple (drawn by A. Drummond)”, where the keyword is “foul”. Therefore, the type of this event is “Foul”. For this semi-structured data, the sentence is parsed into multi-tuples using the character index of the keyword and specific words that appear in the sentence, such as “drawn by”. To incorporate video information into the multi-tuples, the video timestamp is intended to be associated with the text description. OCR is used to identify the time on the scoreboard in each frame and record the timestamp of the current frame. Specifically, we employ the open-source OCR toolkit PaddleOCR [^3] https://github.com/PaddlePaddle/PaddleOCR and Tesseract-OCR [41] to recognize the time simultaneously. By using multiple OCR toolkits, we can improve the overall accuracy of recognition. This is because one toolkit may succeed where the other fails. In addition to event information, player and team information are also stored in multi-tuples. Through these efforts, a Multimodal Basketball Game Knowledge Graph containing 11489 events and 42870 relationships is constructed. Fig .2 shows an example of the Multimodal Basketball Game Knowledge Graph (MbgKG).
TABLE I: Statistics of the events in KG_NBA_2022. “num.” is the abbreviation of “number”. Numbers in bold indicate the highest two values.
num. 986 2649 68 271 1103 1189 774 49 4400
III-B Multimodal Basketball Video Game Captioning Dataset
To construct a dataset that contains videos, text descriptions, and additional knowledge (i.e., player images and names), the pertinent data is extracted from KG_NBA_2022 by utilizing the relations within KG_NBA_2022. In basketball, shot and rebound events are the most prevalent. Meanwhile, we count the number of different types of events in KG_NBA_2022. Table I shows that there are 4400 shot events and 2649 rebound events, which are the most common types of events. Based on the aforementioned common knowledge and statistical data, videos and event descriptions about shot and rebound are extracted from KG_NBA_2022 to construct the dataset. To be more realistic, the text descriptions need to be further modified.
Figure 5: Word cloud of VC_NBA_2022 and Goal datasets. The bigger the font, the more percentage it occupies.
As shown in Fig. 4, the event-action relation in the graph is used to extract 9 kinds of events, and the video-time relation in the graph is used to obtain the video timestamp of the event. Specifically, we first roughly estimate the start and end timestamps of the event based on the existing timestamp. Each clip undergoes a manual review process to ensure accurate start and end timestamps. Subsequently, the video-description relation is used to obtain the text description, which is then matched with the corresponding video clip. We extract player-related information (player images and player names) from KG_NBA_2022 through the team-player and player-name relations. Since each event involves certain players, only the involved individuals and their teammates need to be considered as candidate players for video annotation.
Using the sentence “B. Ingram misses 2-pt jump shot from 19 ft and defensive rebound by J. Winslow” as an example, we will explain our text modification process. Due to the difficulty in generating distance in video captioning task, the “from 19 ft” in the sentence is removed. Based on the attribute information of the player in KG_NBA_2022, the abbreviated name (“J. Winslow”) in the sentence is replaced with the full name (“Justise Winslow”). This change might increase the complexity of name generation but is more aligned with real-world scenarios. To improve text fluency, alterations are made based on the character index. Specifically, we identify the index of keywords like ”defensive” and apply a rule (“defensive/offensive rebound by SOMEONE” $\Longrightarrow$ “SOMEONE gets the defensive/offensive rebound”) to revise the sentence. Finally, there are a total 3977 videos [^4] The scale of the current dataset is not large enough, which is emergent to be enriched in our subsequent research., each of which belongs to one of 9 kinds of events. Each video has one text description and several candidate players information (images and names). We randomly select 3162 clips for training and 786 clips for testing. We also provide word-cloud-based statistics in Fig. 5 (a) to reveal the relative amount of different words. It shows that the top-4 subjects in VC_NBA_2022 are “jump”, “shot”, “3pt”, and “defensive”, followed by “rebound”, “2pt”, “layup”, and “makes”. Table II shows the sample distributions across different labels of events.
In Table III, the comparison of our dataset with MSR-VTT, YouCook and ActivityNet Captions further demonstrates the fine-grained details of our captioning annotations. VC_NBA_2022 has the most sentences per second of 0.182, while the other datasets are all below 0.1, this reflects that our dataset has more detailed information in descriptions. Moreover, VC_NBA_2022 has 1.73 verbs in a sentence on average, higher than 1.41 for ActivityNet Captioning and 1.37 for MSR-VTT. Similarly, the verb ratio of VC_NBA_2022 which is computed by dividing the total number of verbs by the total number of words in the sentence is also much higher than other 3 datasets. This highlights that our dataset is primarily focused on the fine-grained actions of the subjects, aligning with our original intent.
Figure 6: The architecture of knowledge guided entity-aware video captioning network (KEANet) based on a candidate player list in encoder-decoder form.
TABLE II: Statistics of the labels in VC_NBA_2022. “succ.” and “fail.” are abbreviations of “success” and “failure”, respectively.